# SpecTra: Enhancing the Code Translation Ability of Language Models by Generating Multi-Modal Specifications

### v2

@@ Abstract
Large language models (LLMs) are increasingly being used for the task of automated code translation, which has important real-world applications. However, most existing approaches use only the source code of a program as an input to an LLM, and do not consider the different kinds of specifications that can be extracted from a program. In this paper, we propose SpecTra, a multi-stage approach that uses a novel self-consistency filter to first generate high-quality static specifications, test cases, and natural language descriptions from a given program, and then uses these along with the source code to improve the quality of LLM-generated translations. We evaluate SpecTra on three code translation tasks - C to Rust, C to Go, and JavaScript to TypeScript - and show that it can enhance the performance of six popular LLMs on these tasks by up to 10 percentage points and a relative improvement of 26%. Our research suggests that generating high-quality specifications could be a promising and efficient way to improve the performance of LLMs for code translation. We make our code and data available111https://github.com/spectra822/emnlp24, anonymized for review.
SpecTra: Enhancing the Code Translation Ability of Language Models by Generating Multi-Modal Specifications
Vikram Nitin1, Rahul Krishna2, Baishakhi Ray1, 1Columbia University 2IBM T.J. Watson Research Center
1 Introduction
Code translation involves transforming code from one programming language into functionally equivalent code in another language. This task is crucial because new programming languages (e.g., Rust, Go, TypeScript) are continuously developed to overcome the limitations of older ones (e.g., C, JavaScript). Maintaining code written in outdated languages can be costly, susceptible to security vulnerabilities, and challenging to enhance— the “technical debt” associated with maintaining legacy code costs the USA over a trillion dollars annually Mims (2024). Consequently, there is a pressing need to translate legacy code into modern programming languages.
Traditionally, translating code between languages has been accomplished with transpilers Galois (2018); Feldman (1990). These are language-specific rule-based systems that use intricately designed algorithms and heuristics to produce code in the target language that is guaranteed to be functionally identical to the code in the source language. Although technically correct, this code is often non-idiomatic, i.e., very different from typical human-written code. For instance, it may have unusual control flow, use uninformative identifier names, call foreign functions from the original source language, and often contain same problems of the original legacy code. This can be hard to read and maintain, and often defeats the purpose of translating the code in the first place.
Over the past few years, Large Language Models (LLMs) have been employed to tackle a range of code-related tasks, including code generation, code summarization, bug fixing, and test case generation Li et al. (2022); Ahmad et al. (2022); Gao et al. (2023). An expanding body of research Roziere et al. (2020); Ahmad et al. (2022); Roziere et al. (2021); Pan et al. (2024); Yang et al. (2024) focuses on leveraging LLMs for automatic code translation between programming languages. Unlike transpilers, LLMs generate idiomatic code that is more readable, though they offer fewer guarantees of correctness by construction (see Figure 1).
In this paper, we aim to enhance the code translation capabilities of LLMs by integrating the principles of transpilers into the LLM framework. Transpilers are known for their ability to ensure functional correctness, meaning that the translated code maintains the functional specifications of the code written in source language without introducing semantic errors. To leverage this advantage within LLMs, we expose the functional specifications of the source code within the prompt provided to the LLM while translating. By explicitly including these specifications in the prompt, we guide the LLM to adhere to the original functional requirements, thus aiming to combine the readability and idiomatic quality of LLM-generated code with the correctness traditionally associated with transpilers. This method enhances the accuracy and utility of LLMs in code translation tasks, providing a robust solution that balances readability and correctness.
To this end, we propose SpecTra, an approach to enhance the code translation ability of LLMs using multi-modal specifications that capture a program’s functionality. A program’s functionality can be represented in several ways. Static specifications express relationships between inputs and outputs that hold true for all inputs/outputs. Functionality can also be represented through a subset of input/output behavior (I/O specifications), which correspond to specific instances of inputs and outputs to the program (as opposed to all I/Os mentioned in the previous case). Additionally, natural language descriptions, though often less precise and more ambiguous, provide valuable information about a program’s functionality. In the remainder of this paper, we leverage these three modalities—static, dynamic, and natural language specifications—to capture a program’s functionality.
However, inferring specifications of a program is challenging. Traditional methods like Daikon Perkins and Ernst (2004) have limitations: they often fail to scale across different programming languages and the specifications they produce are typically written in some domain specific languages (DSL). These DSLs are not easily understood by LLMs, especially if an LLM has not been the DSL during training. This disconnect poses a significant challenge when trying to utilize LLMs for specification generation. To address this issue, here we leverage a SOTA LLM to generate specs of different modalities of the program written in source language. This approach mirrors using an LLM as an annotator, but with a critical enhancement to ensure the quality and validity of the generated specifications. For I/O spec, we simply run the program to check its validity. For other two types of spec, after the LLM generates the specifications, we ask it to regenerate the original source code based on these specifications. If the regenerated code matches the original implementation, we take this as an indication that the generated specifications are valid and accurately capture the program’s functionality. This validation step is inspired by the self-consistency evaluation method Min et al. (2023) . By ensuring that the LLM can reproduce the original source code from the generated specifications, we increase confidence in the accuracy and usefulness of these specifications. Next, we provide each of these specifications one at a time to an LLM prompt along with the program’s source code, and generate multiple candidate translations in the target language.
@@ 1 Introduction
We evaluate SpecTra on three code translation tasks - converting C to Rust, C to Go, and JavaScript to TypeScript. There is a lot of interest within the software engineering community in converting between these pairs of languages Google (2021a, b), and thus these tasks have direct real-world applicability. We find that SpecTra is able to enhance the performance of 6 popular LLMs on these tasks by up to 26% compared to a uni-modal baseline.
@@
1.
We propose a novel approach based on self-consistency to generate specifications of a program in the form of static specifications, test cases, and natural language descriptions.
2.
We integrate these self-consistent specifications with the source code and propose SpecTra, an approach that utilizes multi-modal specifications to improve the quality of LLM-generated code translations.
3.
We evaluate SpecTra on three program translation tasks and six popular open source and proprietary LLMs of various sizes. The findings of this paper indicate that SpecTra improves the performance of baseline models by up to 26% (relative) and 10% (absolute).
2 Methodology
@@ 2.1.1 Static Specifications
For full example prompts, refer to the supplementary material. The generated static specifications have the following components:
@@
2.1.2 Input-output Specifications
Input-output specifications represent the expected behavior (i.e., output) of a section of a program given a specific input. As with the static specification, input-output specifications () are generated using an LLM for the given program (see Fig. 3).We make use of another prompt composing function () to take a program and create a text prompt containing instructions to generate test inputs () and corresponding outputs (). In other words, . In summary, the input-output specification generation process may be formulated as:
2.1.3 Generating Descriptions
@@ 2.2 Specification validation
We define an evaluator to compare the equivalence (based on an evaluation criteria ) of the variants of the source code () that were generated using with original source code . We may represent this as follows:
Our design allows for a variety of comparator functions to be employed. These may be heuristics that measure code similarity (such as BM25 Robertson and Walker (1994)), or distinct test cases that exercise the program in a specific manner and expect a pre-determined output. This work employs the latter approach in that we used existing test input and expected output pairs (denoted by ) as evaluators. Specifically, our evaluation function would assess if the regenerated code behaves the same as the original source code for all tests.
@@ 2.3 Specification-guided Translation
In this work, we use an ordered sequence of translations starting with the static specification (). If this approach fails, we then use the input-output specifications (), and, if needed, the natural language descriptions (). The sequence of specifications is determined by their formality and completeness, starting with the most structured () and ending with the least formal (). In the rare case that all of these fail, we fall back on the “vanilla” translation with no specifications.
3 Experimental Setup
@@
Datasets. We evaluate our approach on a subset of the CodeNet dataset Puri et al. (2021). CodeNet contains 4053 competitive coding problems, and over 13 million code submissions in a variety of languages including C and JavaScript. These submissions are licensed for public use and redistribution, and anonymized to protect the identity of the authors. We first filter out the problems which don’t have test cases. At random, we select 300 C solutions and 300 JavaScript solutions from the remaining problems, making sure that we don’t pick more than one solution for the same problem. These 300 C and 300 JavaScript programs, each with accompanying test cases, comprise our evaluation dataset.
@@
Models. We evaluate on six large language models:(a) 4 proprietary LLMs: gpt-4o and gpt-3.5-turbo from OpenAI, claude-3-opus from Anthropic; gemini-1.0-pro from Google; and (b) 2 open-source LLMs: Deepseek coder (33B), and Granite (34B). We accessed these models through their respective web APIs. To generate multiple candidate specifications, we use a temperature of 0.6. To generate a single candidate translation, we use greedy decoding with temperature 0, and if we need multiple candidate translations, we use a temperature of 0.3.
@@
Evaluation. We compile each program into an executable, and run the executable with the provided test input. If the output matches the provided test output, then we mark this as correct. If the program doesn’t compile, or exits with an error code, or if the output doesn’t match, then we mark this as incorrect. We create a list of specifications in this order - . If we were unable to generate one modality of specification, then we omit it from the list.
@@
To evaluate translations, we categorize translations into steps. In step 1, we generate translations with just the first specification from the above list and measure the translation accuracy. Next, for step 2, we generate a pair of translations using the first two specifications from the list, and measure translation accuracy for either of the these candidate translations. Lastly, for step 3, we generate using the first three specifications from the list and measure translation accuracy for either of the these candidate translations. In order to establish a baseline, we generate three translations without specifications and measure pass@1 for stage 1, pass@2 for stage 2, and pass@3 for stage 3.
4 Results
4.1 Evaluating the Generated Specifications
We would like to determine whether our specification generation process is able to generate high-quality self-consistent specifications. To do this, we check how many programs in our benchmark dataset we are able to generate self-consistent specifications for. For each program, we generate up to candidate tests, and up to candidate static specifications and descriptions. In all the cases, we stop as soon as we find a self-consistent specification. We use gpt-4o as our LLM for both the “forward” specification generation as well as the “reverse” code generation process.
[…]
The number of self-consistent specifications generated by gpt-4o of each category is shown in Table 1. We also measured the generated test coverage for C programs using the gcov tool. The average coverage was 91.5%, which is an indicator that the tests are of high quality. For the generated descriptions and static specifications, the fact that we are able to recover a functioning program from most of them is a strong validation of their quality. We also manually inspected several of them and confirmed that they are a good representation of the program’s behavior.
[…]

### v3

# SpecTra: Enhancing LLM-based Code Translation with Multi-Modal Specifications

@@ Abstract.
Large language models (LLMs) are increasingly being used for the task of automated code translation, which has important real-world applications. However, most existing approaches use only the source code of a program as an input to an LLM, and do not consider the different kinds of specifications that can be extracted from a program. In this paper, we propose SpecTra, a multi-stage approach that uses a novel self-consistency filter to first generate high-quality static specifications, test cases, and natural language descriptions from a given program, and then uses these along with the source code to improve the quality of LLM-generated translations. We evaluate SpecTra on three code translation tasks - C to Rust, C to Go, and JavaScript to TypeScript - and show that it can enhance the performance of six popular LLMs on these tasks by up to a relative improvement of 46%. We also present a case study on extending this approach to handle translation of a full C project to Rust. Our research suggests that generating high-quality specifications could be a promising and efficient way to improve the performance of LLMs for code translation.
1. Introduction
Code translation refers to the process of converting code written in one programming language into functionally equivalent code in another language. This task is particularly important when updating legacy code written in older languages (e.g., C, JavaScript) to modern programming languages (e.g., Rust, Go, TypeScript). Updating legacy code for critical systems become critical as legacy code often incurs high costs due to maintenance overhead, outdated dependencies, and lack of documentation, making debugging and improvements difficult. It also increases security risks, as older systems lack critical updates and are more vulnerable to modern threats. In fact, the “technical debt" associated with legacy code maintenance costs the United States more than one trillion dollars annually (Mims, 2024). Consequently, translating legacy code into modern languages is an urgent priority.
Traditionally, program translation has been achieved through transpilers (Galois, 2018; Feldman, 1990). Transpilers are language-specific, rule-based systems utilizing carefully designed algorithms and heuristics to produce code in the target language, where the transpiled programs are semantically equivalent (within compilation framework gurantee) to behave identically to the original source code. Although such translations are functionally equivalent, they frequently lack idiomatic structure—meaning that the resulting code differs significantly from typical human-authored programs. For example, transpiler-generated code may exhibit unusual control flow patterns, rely on uninformative or machine-generated identifier names, invoke foreign functions from the source language, and perpetuate the inherent problems of the original legacy code. Such non-idiomatic code can therefore be challenging to read, maintain, and enhance, ultimately undermining the intended benefits of code translation. To address these issues, a growing body of research (Roziere et al., 2020; Ahmad et al., 2022; Roziere et al., 2021; Pan et al., 2024; Yang et al., 2024) investigates the use of LLMs for automatic translation between programming languages. In contrast to transpilers, LLM-generated translations tend to produce code that is human-readable and maintainable, but provides fewer guarantees regarding correctness (see Section˜2).
In this paper, we propose improving LLM-based code translation by integrating core principles from transpiler methodologies. Transpilers rigorously preserve a program’s functional specifications—ensuring that translated code is semantically equivalent to the original. We adopt this key insight by explicitly embedding these functional specifications into the prompts used by the LLM during translation. Rather than relying solely on the LLM’s implicit understanding, we explicitly define functional constraints, guiding the model to produce code that aligns closely with the original semantics. This approach, thus, aims to adapt the semantic equivalency checking from transpiler methods (although at a very approximate level) with the readability and idiomatic nature of LLM-generated code, providing translations that are both accurate and maintainable.
Semantics of a program can be expressed in different ways, such as formal specifications (abbreviated as “specs”) written as logic rules, or natural language (NL) specs. Each helps to ensure correctness and clarity in different ways. While formal specs can define the program’s structure, invariants that the programs must always satisfy, etc., NL specs can explain the program’s overall intent in human-readable form, making it easier to connect the formal rules with human understanding. We hypothesize that, to improve translation accuracy, LLMs benefit from both formal and informal specs, as they provide a deeper understanding of the source program in complementary ways. To this end, we use three different types of specifications:
Static Specifications capture properties of source programs that universally hold for all valid inputs and outputs, such as preconditions, postconditions, and type constraints. These specifications define global correctness criteria independent of particular execution instances. Such specs are often helpful in verification & validation.
Dynamic Specifications illustrate how the program behaves with real input-output examples. Unlike static specs, they give concrete demonstrations of expected functionality in different situations and thus, often help in testing and debugging.
Natural Language Descriptions, although inherently less precise, offer intuitive context and capture informal human insights into a program’s intended functionality. Such descriptions can assist the LLM in understanding the broader semantic goals and idiomatic conventions underlying a given code translation task.
By combining these three types of specifications, we provide structured constraints and intuitive guidance to the LLM during translation. This approach not only improves the translated code while preserving the original semantics, as traditional transpilers do, but also maintains readability, clarity, and natural coding style.
However, inferring precise functional specifications for arbitrary source programs across multiple programming languages is inherently difficult. Traditional specification inference tools (Taghdiri and Jackson, 2007; Ramanathan et al., 2007; Kang et al., 2016), such as Daikon (Perkins and Ernst, 2004), typically focus on specific types of specifications, depend heavily on language-specific instrumentation, and express results using specialized domain-specific languages (DSLs). As a result, these tools may not generalize well across different languages and can be challenging for LLMs to interpret, particularly if the LLMs have not been trained on these DSLs.
To address this limitation, following the previous ML/LLM-based approaches for inferring program properties (Raychev et al., 2015; Rahmani et al., 2021; Endres et al., 2024), we employ a state-of-the-art LLM directly as a multi-modal specification annotator. However, there is no guarantee that the LLM generated specs are actually correct. The key novelty of our approach lies in the rigorous validation of these LLM-generated specifications, ensuring they accurately reflect the semantics of the original source program. Concretely, our validation methodology is tailored specifically to the type of generated specifications:
For static and natural language specifications, we introduce a novel self-consistency-based validation strategy, inspired by recent advancements in robustness testing of code models (Min et al., 2023). Given a source code snippet, we prompt an LLM to generate a specification and then use this specification to regenerate the code in the original language, without any additional context. If the regenerated code closely matches the original (both written in the same source language), it suggests that the specification accurately captures the program’s essential functionality and can be reliably used for translation. Conversely, significant discrepancies indicate gaps or inaccuracies in the generated specification, necessitating further refinement or rejection.
Validation for dynamic specifications (I/O) is straightforward: we execute the original program using the LLM-generated inputs and verify whether the resulting outputs match those generated by the model. Any mismatch indicates incorrect or incomplete specifications. When discrepancies arise, we correct output errors by replacing the LLM-generated outputs with the actual execution results. For other execution-related errors, we iteratively regenerate the I/O examples, retrying until consistent, accurate specifications are obtained.
This iterative validation strategy ensures that only specifications that faithfully encapsulate the source program’s original semantics are subsequently used. After validation, we use these high-quality specifications as explicit constraints in translation prompts to the LLM, guiding it towards generating translations that are both idiomatic and semantically correct. To this end, we introduce SpecTra, a novel approach for specification guided LLM-based code translation. SpecTra provides a recipe on how to combine specs of different modalities for faithful program translation. Thus, by rigorously validating specification quality through self-consistency checks and combining the specs in optimal order, our approach significantly enhances the reliability, robustness, and effectiveness of LLM-based cross-language code translation.
@@ 1. Introduction
We evaluate SpecTra using two settings. First, using self-contained code contest level programs, we perform a thorough empirical analysis of the performance of SpecTra and its various design decisions. Here we evaluate on three code translation tasks - converting C to Rust, C to Go, and JavaScript to TypeScript. There is a lot of interest within the software engineering community in converting between these pairs of languages (Google, 2021a, b), and thus these tasks have direct real-world applicability. We find that SpecTra is able to enhance the performance of 6 popular LLMs on these tasks by up to 46% compared to a baseline without specifications.
To further demonstrate the capabilities of SpecTra, we conduct a case study on full-project translation. While resource constraints limit our focus to C to Rust translation, we rigorously evaluate SpecTra on a benchmark of 24 functions from the cat utility in GNU Coreutils, all covered by system-level test cases. Our results affirm that our specification generation and validation framework seamlessly extends to this challenging setting. Moreover, SpecTra outperforms a no-specification baseline, achieving up to a 10% relative improvement, showcasing its effectiveness in producing high-fidelity translations.
@@
We propose a novel approach based on self-consistency to generate and validate formal and informal specifications of a program.
We demonstrate that strategically combining these specifications enhances program translation quality. To realize this, we introduce SpecTra, a novel framework that seamlessly integrates this approach.
We extensively evaluate SpecTra on three program translation tasks and six popular open source and proprietary LLMs of various sizes.
We further conduct a case study on the full-project translation of a coreutils program(Brady, 2017).
We make our code and data publicly available111https://github.com/spectra822/icse26 , anonymized for review.
| (a) A C program with a complicated conditional, highlighted in yellow. | (b) A static specification for the same C program, generated by an LLM. This also contains an error, highlighted in yellow. |
(f) A correct translation of the C program aided by the above static specification from 1. |
2. Motivating Example
In this section, we introduce an example to illustrate how correct specifications can be beneficial for translating code. This example is taken from CodeNet (Puri et al., 2021), and the responses are produced by GPT-4o. Table˜1 shows a C program which reads an input buffer and prints either "First" or "Second" based on a condition (highlighted).
This condition seems relatively complex on first glance, and it can be tricky to reason about, even for a human. On further analysis, we can see that it evaluates to True if the first and last character of the input are the same and the length of the input is even (suggested by the bitwise AND operation &1.). If we ask an LLM to translate this C code into Rust, as shown in Table˜1, it is able to generate code that reads the input into a buffer and prints "First" or "Second". However, it omits the check for the length of the input being even .
Prior work (Wei et al., 2022) has established that LLMs can benefit from step-by-step prompting on tasks involving complex reasoning. Accordingly, we could ask an LLM to first reason about the properties of the main() function by generating a static specification. However, the first candidate specification that the LLM generates, shown in Table˜1 still has a flaw. This time, the LLM realizes that there is a condition on the length of the input, but it interchanges odd with even. When we try to use this flawed specification to translate the C program, we get an incorrect Rust program shown in Table˜1.
However, if we sample multiple candidate specifications from the LLM by setting a non-zero temperature, we eventually get one that is correct. This is shown at the top of Table˜1. Now, if we provide this correct specification along with the C code, the LLM is successfully able to translate it to Rust, with the correct condition on input length. The resulting translation is shown in Table˜1. Alternatively, we could use dynamic specifications that capture specific inputs and outputs of the program, or natural language specifications that describe the function in imprecise terms. These are also shown in Table˜1. From this discussion, we can derive the following insights:
Insight #1: Correct specifications provided along with code can help an LLM generate a more accurate translation of the code. SpecTra filters out incorrect specifications using a novel self-consistency check, thereby avoiding any potential errors in LLM reasoning. It then uses these self-consistent specifications to improve the quality of LLM translations.
Insight #2: Specifications of multiple modalities can guide the LLM towards a correct translation. SpecTra uses three different modalities of specifications, namely static, dynamic and natural language, all of which work in unison to produce better translations.
3. Methodology
@@
The methodology consists of three steps: (i) specification generation, (ii) specification validation, and (iii) specification combination for translation. This section details each step in depth.
@@ 3.1.1. Static Specifications
Our motivating example, Table˜1, shows a static specification. The generated static specifications have the following components:
@@
3.1.2. Input-output Specifications
Input-output specifications represent the expected behavior (i.e., output) of a section of a program given a specific input. As with the static specification, input-output specifications () are generated using an LLM for the given program. We make use of another prompt composing function () to take a program and create a text prompt containing instructions to generate test inputs () and corresponding outputs (). In other words, . In summary, the input-output specification generation process may be formulated as:
3.1.3. Generating Descriptions
@@ 3.2 Specification validation
We define an evaluator to compare the equivalence of the variants of the source code that were generated using with original source code . We may represent this as follows:
Our design allows for a variety of comparator functions to be employed. These may be heuristics that measure code similarity (such as BM25 (Robertson and Walker, 1994)), or distinct test cases that exercise the program in a specific manner and expect a pre-determined output. This work employs the latter approach in that we used existing test input and expected output pairs (denoted by ) as evaluators. Specifically, our evaluation function would assess if the regenerated code behaves the same as the original source code for all tests.
@@
Handling Multi-Function Standalone Programs. Our static specifications and descriptions are generated per function, i.e., we ask the specification-generation model to annotate each function with a specification. This design choice allows our technique to scale to programs with many functions, where a single specification may not sufficiently express the behavior of the entire program. While verifying self-consistency, we start with a multi-function specification and ask an LLM to generate the entire program.
For I/O specifications, however, we generate only a single specification for an entire program. It is difficult to describe and validate I/O specifications for individual functions, because the arguments and return types can be arbitrarily complex types like nested structures or pointers. Generating program-level test cases sidesteps this difficulty. Further, since these programs are relatively short, the LLM can potentially reason about the flow of information through the program, using the program-level test cases. However, this assumption is no longer valid when we consider large software projects. This is discussed in more detail in Section˜6.1.
@@ 3.3 Specification-guided Translation
In this work, we use an ordered sequence of translations starting with the static specification (). If this approach fails, we then use the input-output specifications (), and, if needed, the natural language descriptions (). The sequence of specifications is determined by their formality and completeness, starting with the most structured () and ending with the least formal (). In the rare case that all of these fail, we fall back on the “vanilla” translation with no specifications. A naive approach is to use all the specification modalities together in the same prompt. We avoided this in favor of a multi-stage approach, guided by the following intuition. As we add each successive specification modality, the information that is gained saturates, but the size of the input keeps increasing. It is known (Liu et al., 2024) that LLMs struggle to process information in long contexts, and this could cause the translation accuracy to degrade.
4. Experimental Setup
4.1 Datasets
@@
Code-Contest Dataset. We evaluate our approach on a subset of the CodeNet dataset (Puri et al., 2021). CodeNet contains 4053 competitive coding problems, and over 13 million code submissions in a variety of languages including C and JavaScript. We first filter out the problems which don’t have test cases. At random, we select 300 C solutions and 300 JavaScript solutions from the remaining problems, making sure that we don’t pick more than one solution for the same problem. These 300 C and 300 JavaScript programs, each with accompanying test cases, comprise our evaluation dataset.
Full-project Translation Case Study. We use the cat utility from GNU Coreutils222https://github.com/coreutils/coreutils. It consists of over a hundred functions, but several of these are in header files that are shared among multiple coreutils utilities. We pick a subset of 24 of these functions, that are covered by the end-to-end test cases provided for cat in the Coreutils repository. They have 1452 lines of code, with an average of 60.5 lines per function. The longest function is 522 lines long, and the shortest function is 5 lines long.
We write custom LLVM/Clang (Lattner and Adve, 2004) compiler passes to decompose the entire C program and extract the source code and docstrings of individual C functions. We also instrument the code by adding an LLVM pass that inserts fprintf statements at the end of each function body, just before the return statement, to log the arguments and returned value. Printing structures and pointers needs special care. We iterate through individual fields of structures and log their values. For pointers, if it is a char* pointer, then we treat it as a string and print up to its first 100 characters. If it is any other pointer type, then we simply print its address.
Once we translate a function from C to Rust, we need to make sure that the code runs end-to-end. To do this, we use Rust’s C FFI (Foreign Function Interface) to call functions in C from Rust and vice-versa. We use bindgen333https://github.com/rust-lang/rust-bindgen to generate bindings for each function in C. For more details, refer to our code implementation.
4.2 Models
@@
We evaluate on six large language models: (a) 4 proprietary LLMs: gpt-4o and gpt-4o-mini from OpenAI, claude-3-sonnet-v2 from Anthropic; gemini-1.5-pro from Google; and (b) 2 open-source LLMs: Mistral Large 2 (123B), and Llama3.1 (70B). We accessed these models through their OpenAI or VertexAI APIs. For the case study on full-project translation, we use only GPT-4o. To generate multiple candidate specifications, we use a temperature of 0.6. To generate a single candidate translation, we use greedy decoding with temperature 0, and if we need multiple candidate translations, we use a temperature of 0.3.
4.3 Evaluation
@@
To evaluate a translated program, we compile it, and run the executable with the provided ground truth test input. If the output matches the ground truth test output, then we mark the translation as correct. If the program doesn’t compile, exits with an error code, or if the output doesn’t match, then we mark this as incorrect.
@@
For running SpecTra on a program, we generate a list of specifications of different modalities in this order - . If we were unable to generate one modality of specification, then we omit it from the list. We generate 3 candidate translations, using 3 steps. In step 1, we generate translations with just the first specification from the above list and check whether the translation is correct. Next, for step 2, we generate a pair of translations using the first two specifications from the list, and check whether either of them is correct. Lastly, for step 3, we translate using the first three specifications from the list and check whether any of them is correct. Then, pass@1, pass@2 and pass@3 for SpecTra correspond to the accuracy in step 1, step 2 and step 3 respectively. In order to establish a baseline, we generate three translations without specifications and similarly measure pass@1 as the accuracy of just the first translation, pass@2 as the accuracy of the first or the second, and so on.
To this end, we seek to answer the following research questions, evaluating various design choices of SpecTra on smaller standalone programs (see Section˜5).
RQ1: SpecTra Performance with Multi-modal Specs. What is the overall performance of SpecTra?
RQ2: Impact of Individual Specification Modalities. How does each individual specification modality guide the LLM towards better translations?
RQ3: Design Choices. How do the different components of our design, such as the multi-stage pipeline and the self-consistency check, contribute to the overall performance?
RQ4: Evaluating the Generated Specifications. What is the quality of the generated specifications?
Once we know the right recipe to configure SpecTra, we perform a case study to evaluate whether SpecTra can be effectively extend to translate full software projects, as discussed in Section 6.
5. Empirical Results
Here, we measure SpecTra on smaller standalone programs. This helps us to find out the optimal configuration of SpecTra that can be applied to larger real-world setting.
| C to Rust | |||||||
| No. correct | improvement % | ||||||